OpenStudy (anonymous):

f(4)=1 f ' (x)=-2 Find an equation of the tangent line to f(x)at the point x=4.

OpenStudy (anonymous):

to represent a line by algebraic forms, you need to know two things. slope and the point it passes. those informations are in the f(4)=1 f'(x)=-2

OpenStudy (anonymous):

I have to find the slope and y intercept so I don't have any of that either

OpenStudy (anonymous):

I have no idea where to start I really need help please anyone

OpenStudy (leksi):

y-f(4)=f'(x)(x-4) y-1=-2(x-4) 2 x+y-9=0

OpenStudy (anonymous):

so then the 2x+y=0 is the tangent equation

OpenStudy (anonymous):

oh sorry I mean 2x+y-9=0 is the tangent equation

OpenStudy (anonymous):

is this correct

OpenStudy (anonymous):

you understand that f'(a) is a slope of the tangent line at x=a right?

OpenStudy (anonymous):

and f'(x)=-2 says f'(4)=-2

OpenStudy (anonymous):

also, y intercept is just one of the points that line passes. you can use other points that line pass. f(4)=1 says the tangent line passes (4,1)

OpenStudy (anonymous):

so assuming that you understand y=m(x-a)+b, y=(-2)(x-4)+1

OpenStudy (anonymous):

or in other form, 2x+y=9

OpenStudy (anonymous):

ok so then are yousaying that the slope is -2 and the y intercept is 1

OpenStudy (anonymous):

defintion of y intercept is the value of y at x=0

OpenStudy (anonymous):

y intercept is of course 9..

OpenStudy (anonymous):

ok yup just got that so then the slope is -2

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok got that then what it I have t estimate f(7) then what so I do

OpenStudy (anonymous):

i don't understand what you want to know. what is t?

OpenStudy (anonymous):

sorry bad typing I meant how do I find the estimate of f(7)

OpenStudy (anonymous):

oh so estimation is something close to the real solution right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

in this case, (assuming you don't know integral calculus) you can't figure out the exact f(7) because you don't know the rule

OpenStudy (anonymous):

right it just has to be an estimate

OpenStudy (anonymous):

yeah. well you can use your tangent line to make that approximation

OpenStudy (anonymous):

ok so how do you do that

OpenStudy (anonymous):

you plug 7 into x in your equation of tangent line

OpenStudy (anonymous):

actually the whole point of calculus is that if you cut curves very small, it estimates to straight lines

OpenStudy (anonymous):

so then is the esimate of f(7)=-5

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

it is called 'linear approximation' and your answer is right

OpenStudy (anonymous):

yay ok thanks for all your help